# Hydrostatic mass of galaxy clusters within some theories of gravity

### v2

@@ 4 Concluding remarks
In this work, we have revisited the derivation of MENG with GUP correction derived in Ref. [17]. We also derive the hydrostatic mass of GCs within EiBI theory and BHG. We applied all formulations that we have revisited and derived on 10 GCs and compared the results with the Newtonian mass obtained from Ref. [3]. The impact of the theories of gravity used in this work starts if we set $\beta_{0}=-1.656\times 10^{110}$ for MENG, $\kappa=5\times 10^{40}$ for EiBI theory, and $\Upsilon=-0.1655\times 10^{69}$ for BHG. If the values of the free parameters are set closer to zero than the values mentioned before, the impact is negligible, which means that one can safely use the Newtonian formulation to calculate the hydrostatic mass of GCs.
We compare our results with $M_{bar}$. The best linear fit is achieved by MENG with $\beta_{0}=2.70\times 10^{115}$, which gives the slope $\mathcal{M}$ of $0.755\pm 0.061$. This is the finest value that we found to reach the finest slope that is closest to unity. The linear fits of EiBI theory and BHG theory show much lower values of the slope compared to the MENG’s slope value. This leads us to the conclusion that in this study, MENG is the best theory of gravity that alleviates the mass discrepancy of hydrostatic mass and baryonic mass of GCs.

### v3

@@ 4 Concluding remarks
In this work, we have revisited the derivation of MENG with GUP correction derived in Ref. [19], and derive the hydrostatic mass of GCs within MENG with GUP correction. We also derive the hydrostatic mass of GCs within EiBI theory and BHG. We applied all formulations that we have revisited and derived on 10 GCs and compared the results with the Newtonian mass obtained from Ref. [3]. The impact of the theories of gravity used in this work starts if we set $\beta_{0}=-1.656\times 10^{110}$ for MENG, $\kappa=5\times 10^{40}$ for EiBI theory, and $\Upsilon=-0.1655\times 10^{69}$ for BHG. If the values of the free parameters are set closer to zero than the values mentioned before, the impact is negligible, which means that one can safely use the Newtonian formulation to calculate the hydrostatic mass of GCs.
We compare our results of EiBI theory and BHG with $M_{bar}$. A better linear fit is achieved by EiBI theory with $\kappa_{0}=5.80\times 10^{40}$, which gives the slope $\mathcal{M}$ of $0.126\pm 0.086$. This value is closer to unity than the one of BHG. This leads us to the fact that EiBI theory is more effective than BHG in alleviating the mass discrepancy between hydrostatic mass and baryonic mass in GCs. Nevertheless, neither EiBI theory nor BHG completely addresses the mass discrepancy problem.